# CT-AGRG: Automated Abnormality-Guided Report Generation from 3D Chest CT Volumes

### v2

@@
The rapid increase of computed tomography (CT) have created a need for robust automated analysis techniques in clinical settings to assist radiologists managing their growing workload. Existing methods generate entire reports directly from 3D CT images, without explicitly focusing on observed abnormalities. This unguided approach can result in repetitive content or incomplete reports. We propose a new anomaly-guided report generation model, which first predicts abnormalities and then generates targeted descriptions for each. Evaluation on a public dataset demonstrates significant improvements in report quality and clinical relevance. We extend our work by conducting an ablation study to demonstrate its effectiveness.
@@ 1 Introduction
Three-dimensional Computed Tomography (3D CT) scans have emerged as indispensable tools in medical imaging [1]. Given the rapidly increasing number of scans to analyze [2, 3, 4], the time-consuming nature of the task [5] and the important demand for specialized radiological expertise in numerous healthcare systems [6, 7], computed-aided diagnosis and automating report generation have emerged as an active research area  [8, 9, 10]. Whether dealing with 2D [11, 12] or 3D [13, 14] images, numerous deep learning methods have been introduced to support healthcare professionals in tasks such as anomaly classification [13, 15], detection [16, 17, 18], segmentation [19, 20, 21], and report generation [9, 10, 22, 23]. Indeed, the availability of public datasets for 2D radiology images [12, 11] has facilitated the development of such methods [22, 23, 24, 25, 26, 27]. Recently, CT2Rep [10] has been presented as the first 3D method for automated report generation, trained and evaluated on the public CT-RATE dataset [14] of CT examination. CT2Rep, uses an Encoder-Decoder model to generate reports directly from 3D CT scans. However, this end-to-end approach may lead to incomplete [6]. Reading a CT study and composing a report [28], a radiologist detects abnormalities and then provide descriptions for each identified abnormality. Inspired by the radiologists’ workflow, we propose a method that first predicts anomalies and then generates a specific sentence for each anomaly. Our contributions are as follows:
@@
  • We leverage a pre-trained language model that we condition to generate a sentence for each abnormality.
@@ 4.1 Multi-label classification pre-training
The model receives a volume $x\in\mathbb{R}^{240\times 480\times 480}$ as input. This volume is passed to a visual feature extractor $\Phi_{V}$, either CT-Net [13] or CT-ViT [47]. CT-Net consists of ResNet [54] modules pre-trained on ImageNet [55] followed by a 3D convolutional network [56]. CT-ViT [47] is a Vision Transformer [40] based on the attention mechanism [41] computed from 3D patches extracted from the initial volume. From the initial volume $x$, both CT-Net and CT-ViT yield a vector representation $h\in\mathbb{R}^{512}$, such that:
@@ 4.2 Step 1: Multi-task single label classification
The multi-label classification head $\Psi$ of the pre-trained network is replaced by 18 independent projection heads [59], denoted as $\Psi^{p}_{i}$ ($i\in\{1,\ldots,18\}$), followed by classification heads denoted as $\Psi^{c}_{i}$ ($i\in\{1,\ldots,18\}$). The feature vector $h$ is first passed through each projection head $\Psi^{p}_{i}$ to obtain a unique embedding noted $h_{i}$ per class label, such that:
@@ 4.3 Step 2: Abnormality-based sentence generation
Method BL-1 BL-4 M RL BERTs BARTs P R F1
CT2Rep [10] 0.323 0.115 0.176 0.241 0.864 -3.178 0.416 0.367 0.360
CT-AGRG w/ CT-ViT 0.330 0.137 0.209 0.301 0.866 -2.962 0.349 0.638 0.426
CT-AGRG w/ CT-Net 0.392 0.175 0.209 0.318 0.871 -2.900 0.443 0.650 0.515
@@
BL-4 M P R F1
0.083 0.127 0.404 0.190 0.208
0.171 0.194 0.429 0.617 0.492
0.151 0.197 0.430 0.640 0.503
0.175 0.209 0.443 0.650 0.515
Table 1: Quantitative evaluation on the test set with Natural Language Generation (NLG) metrics and Clinical Efficacy (CE) metrics. Our model outperforms the previous method both on NLG and CE metrics. Ablation study: The evaluation on NLG and CE metrics highlights the contribution of each module of our method. The ablation study on CT-AGRG was conducted with CT-Net [13] as visual feature extractor, since it yields the best results.
@@ 6.1 Quantitative results
Abnormalities mentioned in the generated report are extracted with the RadBERT labeler [14, 52]. These predicted anomalies are compared against ground truth labels to compute Clinical Efficacy (CE) metrics: Precision, Recall, and F1-Score [79]. We also evaluate our captioning model using widely used [10, 22, 23, 26] Natural Language Generation (NLG) metrics: BLEU-1 and BLEU-4 [80] for n-gram precision; METEOR [81] for alignment with stemming and synonymy; ROUGE [82] for overlap focusing on recall; BERT-Score [83] for token-level contextual embedding similarity; and BART-Score [84] for sequence-level likelihood assessment. This combination captures lexical, semantic, and generative aspects of caption quality, providing a comprehensive evaluation of our model’s performance [85, 86]. Table 1 demonstrates that CT-AGRG achieves significantly better results than prior work. With CT-ViT and CT-Net as visual feature extractors, we achieve an average Recall of 0.638 and 0.650 ($\Delta$+74and $\Delta$+77increase compared to CT2Rep [10]), and an average F1-Score of 0.426 and 0.515, respectively ($\Delta$+18and $\Delta$+43increase). CT-AGRG achieves a BARTScore [84] of -2.900, outperforming CT2Rep [10] (-3.178, $\Delta$+9increase), indicating enhanced abnormality detection capabilities and improved semantic relevance in generated captions.
@@ 6.2 Ablation study
As shown in Table 1, we conduct an ablation study to evaluate each component of our CT-AGRG model. Introducing multi-task classification significantly improves F1-Score from 0.208 to 0.492 ($\Delta$+136increase). Augmenting the latent space further enhances performance, increasing Recall from 0.617 to 0.650 ($\Delta$+5increase) and F1-Score from 0.492 to 0.515 ($\Delta$+5increase). NLG metrics show notable improvements, with BLEU-4 [80] and METEOR [81] scores increasing by 0.092 and 0.082, respectively, indicating enhanced sentence variability while maintaining semantic fidelity to ground truth reports.
@@ 8 Acknowledgments
Appendix A Quantitative results
CT2Rep CT-AGRG (Ours) Abnormality P R F1 P R F1 Frequency Medical material 0.461 0.120 0.190 0.630 0.443 0.521 0.114 Arterial wall calcification 0.679 0.751 0.713 0.681 0.872 0.765 0.267 Cardiomegaly 0.476 0.422 0.447 0.488 0.774 0.599 0.108 Pericardial effusion 0.136 0.085 0.105 0.269 0.505 0.351 0.074 Coronary artery wall calcification 0.690 0.701 0.696 0.660 0.880 0.755 0.244 Hiatal hernia 0.186 0.510 0.273 0.217 0.701 0.331 0.134 Lymphadenopathy 0.345 0.526 0.417 0.397 0.646 0.491 0.266 Emphysema 0.311 0.479 0.377 0.423 0.431 0.427 0.195 Atelectasis 0.348 0.267 0.302 0.356 0.610 0.450 0.232 Lung nodule 0.547 0.732 0.626 0.622 0.853 0.719 0.425 Lung opacity 0.841 0.431 0.570 0.680 0.787 0.729 0.374 Pulmonary fibrotic sequela 0.358 0.317 0.337 0.393 0.724 0.509 0.267 Pleural effusion 0.716 0.439 0.544 0.639 0.800 0.710 0.126 Mosaic attenuation pattern 0.172 0.315 0.223 0.271 0.533 0.359 0.078 Peribronchial thickening 0.213 0.122 0.155 0.233 0.589 0.334 0.098 Consolidation 0.456 0.303 0.364 0.459 0.653 0.539 0.172 Bronchiectasis 0.357 0.032 0.058 0.294 0.468 0.362 0.099 Interlobular septal thickening 0.200 0.053 0.084 0.260 0.426 0.323 0.071 Mean 0.416 0.367 0.360 0.443 0.650 0.515 0.186
Table 2: Quantitative evaluation on the test set with Precision (P), Recall (R) and F1-Score (F1) for each pathology. Frequency of each pathology in the test set is also provided. The metrics associated with CT-AGRG were computed using CT-Net [13] as the visual feature extractor, as this configuration yielded the best performance. The underlined metrics are those that have improved with our contribution compared to the previous method. Our method improves the classification metrics for the majority of abnormalities.
Appendix B Qualitative results
Fig. 2: Comparison of ground truth with generated reports with CT2Rep [10] and CT-AGRG on test set from CT-RATE [14].
Appendix C Training steps
Fig. 3: Overview of the method with details for each step.
Appendix D Hyperparameters
Step 1
w/ CT-Net [13]
Step 1
w/ CT-ViT [47]
Step 2
w/ CT-Net [13]
Step 2
w/ CT-ViT [47]
Optimizer Adam [69] Adam [69] Adam [69] Adam [69]
Learning rate - Visual Extractor 0.0001 0.0001 0.00001 0.00001
Learning rate - Classification Head(s) 0.0001 0.0001 0.00001 0.00001
Batch size 4 4 4 4
Epochs 50 50 2 2
Early stopping
Table 3: Hyperparameters of the training for each model on the classification steps 1 and 2. For each configuration, we use the hyper-parameters that yield the best results. Early stopping is performed to obtain the checkpoint with the highest F1-Score on the validation set.
CT2Rep [10] CT-AGRG
Optimizer AdamW [70] AdamW [70]
Learning rate - Visual Extractor 0.00005 -
Learning rate - Encoder Decoder 0.00005 0.00005
Weight decay 0.01 0.01
Batch size 16 64
Epochs 100 100
Early stopping
Tokenizer - Max. sequence length 400 / report 60 / sentence
Generation mode Beam search Beam search
Beam size 4 4
Table 4: Hyperparameters of the training for each model on the generation step. For each configuration, we use the hyper-parameters that yield the best results. Early stopping is performed to obtain the checkpoint with the highest F1-Score on the validation set, triggered only after a minimum of 10 epochs to ensure sufficient training time and comparable iterations across experiments.
Appendix E Dataset statistics
Fig. 4: Frequency of abnormalities in train, validation and test sets.
Fig. 5: Token count distribution per sentence from reports for train, validation and test sets.
Fig. 6: Manufacturer distribution for train, validation and test sets.
Fig. 7: Age distribution for train, validation and test sets.
Fig. 8: Token count distribution from reports for train, validation and test sets.
Fig. 9: Token count distribution per sentence from reports for train, validation and test sets.

### v6

@@
The rapid increase of Computed Tomography examinations have created a need for robust automated analysis techniques in clinical settings to assist radiologists managing their growing workload. Existing methods generate entire reports directly from 3D CT images, without explicitly focusing on observed abnormalities. This unguided approach can result in repetitive content or incomplete reports. We propose a new anomaly-guided report generation model, which first predicts abnormalities and then generates targeted descriptions for each. Evaluation on a public dataset demonstrates significant improvements in report quality and clinical relevance. We extend our work by conducting an ablation study to demonstrate its effectiveness.
@@ 1 Introduction
Three-dimensional Computed Tomography (3D CT) scans have emerged as indispensable tools in medical imaging [1]. Given the rapidly increasing number of scans to analyze [2, 3, 4], the time-consuming nature of the task [5] and the important demand for specialized radiological expertise in numerous healthcare systems [6, 7], computed-aided diagnosis and automating report generation have emerged as an active research area  [8, 9, 10]. Whether dealing with 2D [11, 12] or 3D [13, 14] images, numerous deep learning methods have been introduced to support healthcare professionals in tasks such as anomaly classification [13, 15], detection [16, 17, 18], segmentation [19, 20, 21], and report generation [9, 10, 22, 23]. Indeed, the availability of public datasets for 2D radiology images [12, 11] has facilitated the development of such methods [22, 23, 24, 25, 26, 27]. Recently, CT2Rep [10] has been presented as the first 3D method for automated report generation, trained and evaluated on the public CT-RATE dataset [14] of CT examination. CT2Rep, uses an Encoder-Decoder model to generate reports directly from 3D CT scans. However, this end-to-end approach may lead to incomplete reports [6]. Reading a CT study and composing a report [28], a radiologist detects abnormalities and then provides descriptions for each identified abnormality. Inspired by the radiologists’ workflow, we propose a method that first predicts anomalies and then generates a specific sentence for each anomaly. Our contributions are as follows:
@@
  • We leverage a pre-trained language model with biomedical domain knowledge that we condition to generate a sentence for each abnormality.
@@ 4.1 Multi-label classification pre-training
The model receives a volume $x\in\mathbb{R}^{240\times 480\times 480}$ as input. This volume is passed to a visual feature extractor $\Phi_{V}$, either CT-Net [13] or CT-ViT [47]. CT-Net consists of ResNet [54] modules pre-trained on ImageNet [55] followed by a 3D convolutional network [56]. CT-ViT [47] is a Vision Transformer [40] based on the attention mechanism [41] computed from 3D patches extracted from the initial volume. From the initial volume $x$, both CT-Net and CT-ViT yield a vector representation $h\in\mathbb{R}^{2048}$, such that:
@@ 4.2 Step 1: Multi-task single label classification
The multi-label classification head $\Psi$ of the pre-trained network is replaced by 18 independent projection heads [59], denoted as $\Psi^{p}_{i}$ ($i\in\{1,\ldots,18\}$), followed by classification heads denoted as $\Psi^{c}_{i}$ ($i\in\{1,\ldots,18\}$). The feature vector $h$ is first passed through each projection head $\Psi^{p}_{i}$ to obtain a unique embedding noted $h_{i}\in\mathbb{R}^{1024}$ per class label, such that:
@@ 4.3 Step 2: Abnormality-based sentence generation
Method METEOR ROUGEL BERT BART P R F1
CT2Rep [10] $0.173\text{\scriptsize$\pm 0.002$}$ $0.243\text{\scriptsize$\pm 0.002$}$ $0.865\text{\scriptsize$\pm 0.001$}$ $-3.314\text{\scriptsize$\pm 0.085$}$ $41.63\text{\scriptsize$\pm 2.01$}$ $38.12\text{\scriptsize$\pm 0.88$}$ $36.73\text{\scriptsize$\pm 0.76$}$
CT-AGRG w/ CT-ViT $0.190\text{\scriptsize$\pm 0.004$}$ $0.270\text{\scriptsize$\pm 0.002$}$ $0.864\text{\scriptsize$\pm 0.001$}$ $-2.936\text{\scriptsize$\pm 0.040$}$ $37.75\text{\scriptsize$\pm 0.46$}$ $55.35\text{\scriptsize$\pm 0.91$}$ $42.13\text{\scriptsize$\pm 0.23$}$
CT-AGRG w/ CT-Net $\mathbf{0.196}\text{\scriptsize$\pm 0.008$}$ $\mathbf{0.280}\text{\scriptsize$\pm 0.003$}$ $\mathbf{0.867}\text{\scriptsize$\pm 0.001$}$ $\mathbf{-2.714}\text{\scriptsize$\pm 0.179$}$ $\mathbf{45.74}\text{\scriptsize$\pm 0.35$}$ $\mathbf{62.59}\text{\scriptsize$\pm 0.14$}$ $\mathbf{50.11}\text{\scriptsize$\pm 0.15$}$
@@
BLEU-4 METEOR R F1
$0.083\text{\scriptsize$\pm 0.003$}$ $0.127\text{\scriptsize$\pm 0.008$}$ $19.05\text{\scriptsize$\pm 0.19$}$ $23.80\text{\scriptsize$\pm 0.10$}$
$0.154\text{\scriptsize$\pm 0.008$}$ $0.186\text{\scriptsize$\pm 0.005$}$ $56.56\text{\scriptsize$\pm 0.12$}$ $46.11\text{\scriptsize$\pm 0.13$}$
$0.146\text{\scriptsize$\pm 0.005$}$ $0.189\text{\scriptsize$\pm 0.007$}$ $58.89\text{\scriptsize$\pm 0.12$}$ $47.22\text{\scriptsize$\pm 0.11$}$
$\mathbf{0.172}\text{\scriptsize$\pm 0.007$}$ $\mathbf{0.196}\text{\scriptsize$\pm 0.008$}$ $\mathbf{62.59}\text{\scriptsize$\pm 0.14$}$ $\mathbf{50.11}\text{\scriptsize$\pm 0.15$}$
Table 1: Quantitative evaluation on the test set with Natural Language Generation (NLG) metrics and Clinical Efficacy (CE) metrics. Reported mean and standard deviation metrics were computed over five independent runs. Our model outperforms the previous method both on NLG and CE metrics. Ablation study: The ablation study on CT-AGRG was conducted with CT-Net [13] as visual feature extractor, since it yields the best results. It highlights the importance of each module.
@@ 6.1 Quantitative results
Abnormalities mentioned in the generated report are extracted with the RadBERT labeler [14, 52]. These predicted anomalies are compared against ground truth labels to compute Clinical Efficacy (CE) metrics: Precision, Recall, and F1-Score [79]. We also evaluate our captioning model using widely used [10, 22, 23, 26] Natural Language Generation (NLG) metrics: BLEU-4 [80] for n-gram precision; METEOR [81] for alignment with stemming and synonymy; ROUGE [82] for overlap focusing on recall; BERT-Score [83] for token-level contextual embedding similarity; and BART-Score [84] for sequence-level likelihood assessment. This combination captures lexical, semantic, and generative aspects of caption quality, providing a comprehensive evaluation of our model’s performance [85, 86]. Table 1 demonstrates that CT-AGRG achieves significantly better results than prior work. With CT-ViT and CT-Net as visual feature extractors, we achieve an average Recall of 55.35 and 62.59 ($\Delta$+45.20and $\Delta$+64.19increase compared to CT2Rep [10]), and an average F1-Score of 42.13 and 50.11, respectively ($\Delta$+14.70and $\Delta$+36.43increase). CT-AGRG achieves a BART-Score [84] of -2.714, outperforming CT2Rep [10] (-3.314, $\Delta$+18.11increase), indicating enhanced abnormality detection capabilities and improved semantic relevance in generated captions.
@@ 6.2 Ablation study
As shown in Table 1, we conduct an ablation study to evaluate each component of our CT-AGRG model. Introducing multi-task classification significantly improves F1-Score from 23.80 to 46.11 ($\Delta$+93.76increase). Augmenting the latent space further enhances performance, increasing Recall from 56.56 to 58.89 ($\Delta$+4.12increase) and F1-Score from 46.11 to 47.22 ($\Delta$+2.41increase). NLG metrics show notable improvements, with BLEU-4 [80] and METEOR [81] scores increasing by 0.089 and 0.069, respectively, indicating enhanced sentence variability while maintaining semantic fidelity to ground truth reports.